What is the magnitude of the electric force between a proton and an electron when they are at a distance of 3.06 angstrom from each other? Submit Answer Tries 0/12 What is the electric potential produced by the proton at this distance from the proton? Submit Answer Tries 0/12 What is the potential energy of the proton-electron system when they are at this same distance from each other? Submit Answer Tries 0/12what is the magnitude between an electron and proton...

I need to figure out all three questions, thanks!
Answer #1

from Couloumbs law of electro static forces

force between two charges is given by F = Kq1q2/r^2


here q1 = charge of proton = 1.6 *10^-19 C

q2 = charge of electron = -1.6 *10^-19 C

r is the distance between the charges = 3.06 Angstrom units

1 Angstrom unit = 10^-10 m

k is permittivity Constant = 9*10^9

so

F = (9*10^9* 1.6 *10^-19 * 1.6 *10^-19)/(3.06 *10^-10 * 3.06 *10^-10)

F = 2.46 *10^-19 N

------------------------------------

electric potential by proton = V = Kq1/r

V = (9*10^9 * 1.6*10^-19)/(3.06 *10^-10)

V = 4.7 V

-------------------------------------

Elecric potential energy U = Kq1q2/r

U = (9*10^9 * 1.6 *10^-19*1.6 *10^-19)/( 3.06 *10^-10)

U = 7.529 *10^-19 Joules